# On discrete Sobolev inequalities for nonconforming finite elements under a semi-regular mesh condition

### v1

@@ 4 Lemmata for analysis
Furthermore, the operator $\mathcal{L}$ maps $\mathcal{D}(D)\cap L^{1}_{0}(D)$ into $\mathcal{D}(D)^{d}$ and for each real $r\in(1,\infty)$, is continuous from $L^{r}_{0}(D)$ into $W^{1,r}_{0}(D)^{d}$; there exists a constant $C_{r}$ such that
$\displaystyle|\mathcal{L}f|_{W^{1,r}(D)^{d}}\leq C_{r}\|f\|_{L^{r}(D)}\quad\forall f\in L^{r}_{0}(D),$ (4.2)
where $L^{r}_{0}(D):=\{v\in L^{r}(D):\ \int_{D}vdx=0\}$.
@@
A proof is available in [5]; see also [1, Theorem 1.4.5] for further reference. ∎
@@
A proof can be found in [7, Lemma 3]. ∎
@@
Let $p^{\prime}\in[1,\infty)$. For any $\bm{w}\in W^{1,p^{\prime}}(\Omega)^{d}$, we set $\bm{\tau}_{h}:={I}_{h}^{RT}\bm{w}$ in (4.3). Using the definition of $I_{h}^{RT}$ yields (4.4). ∎
@@
Let $p^{\prime}\in(1,\infty)$ and $h\leq 1$. For any $\bm{w}\in W^{1,p^{\prime}}(\Omega)^{d}$ and $\psi_{h}\in V_{h}^{DC(1)}$,
$\displaystyle\left|\sum_{F\in\mathcal{F}_{h}^{i}}\int_{F}\{\!\{\bm{w}\}\!\}_{\omega,F}\cdot\bm{n}_{F}\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}ds\right|$ $\leq c|\psi_{h}|_{p,J}\|\bm{w}\|_{W^{1,p^{\prime}}(\Omega)^{d}},$ (4.5)
$\left|\sum_{F\in\mathcal{F}_{h}^{\partial}}\int_{F}(\bm{w}\cdot\bm{n}_{F})\Pi_{F}^{0}\psi_{h}ds\right|$ $\leq c|\psi_{h}|_{p,J}\|\bm{w}\|_{W^{1,p^{\prime}}(\Omega)^{d}},$ (4.6)
where $\frac{1}{p}+\frac{1}{p^{\prime}}=1$.
@@
Suppose that $F\in\mathcal{F}_{h}^{i}$ with $F=T_{+}\cap T_{-}$, $T_{+},T_{-}\in\mathbb{T}_{h}$. Using the Hölder’s inequality, the weighted average and the trace inequality (2.1) yields
@@
$\quad\leq c\left(\|\bm{w}_{+}\|_{L^{p^{\prime}}(T_{+})^{d}}+h_{T_{+}}^{\frac{1}{p^{\prime}}}\|\bm{w}_{+}\|_{L^{p^{\prime}}(T_{+})^{d}}^{1-\frac{1}{p^{\prime}}}|\bm{w}_{+}|_{W^{1,p^{\prime}}(T_{+})^{d}}^{\frac{1}{p^{\prime}}}\right)\omega_{T_{+},F}\ell_{T_{+},F}^{-\frac{1}{p^{\prime}}}\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}$
$\quad\quad+c\left(\|\bm{w}_{-}\|_{L^{p^{\prime}}(T_{-})^{d}}+h_{T_{-}}^{\frac{1}{p^{\prime}}}\|\bm{w}_{-}\|_{L^{p^{\prime}}(T_{-})^{d}}^{1-\frac{1}{p^{\prime}}}|\bm{w}_{-}|_{W^{1,p^{\prime}}(T_{-})^{d}}^{\frac{1}{p^{\prime}}}\right)\omega_{T_{-},F}\ell_{T_{-},F}^{-\frac{1}{p^{\prime}}}\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}$
$\quad\leq c\Bigg{\{}\left(\|\bm{w}_{+}\|_{L^{p^{\prime}}(T_{+})^{d}}+h_{T_{+}}^{\frac{1}{p^{\prime}}}\|\bm{w}_{+}\|_{L^{p^{\prime}}(T_{+})^{d}}^{1-\frac{1}{p^{\prime}}}|\bm{w}_{+}|_{W^{1,p^{\prime}}(T_{+})^{d}}^{\frac{1}{p^{\prime}}}\right)^{p^{\prime}}$
@@
Using the Hölder’s inequality again, we have
@@
$\quad\leq c\sum_{F\in\mathcal{F}_{h}^{i}}\left(\omega_{T_{+},F}^{p}\ell_{T_{+},F}^{-\frac{p}{p^{\prime}}}+\omega_{T_{-},F}^{p}\ell_{T_{-},F}^{-\frac{p}{p^{\prime}}}\right)^{\frac{1}{p}}\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}$
@@
$\quad\leq c\left(\sum_{F\in\mathcal{F}_{h}^{i}}\left(\omega_{T_{+},F}^{p}\ell_{T_{+},F}^{-\frac{p}{p^{\prime}}}+\omega_{T_{-},F}^{p}\ell_{T_{-},F}^{-\frac{p}{p^{\prime}}}\right)\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}^{p}\right)^{\frac{1}{p}}$
@@
$\quad\leq c|\psi_{h}|_{p,J}\|\bm{w}\|_{W^{1,p^{\prime}}(\Omega)^{d}},$
which is the inequality (4.5) together with the weight (2.9) and Jensen inequality. Here, $\mathbb{T}_{F}$ denotes the set of the simplices in $\mathbb{T}_{h}$ that share $F$ as a common face.
By an analogous argument, the estimate (4.6) holds. ∎
@@ 5 Main theorem
Let $h\leq 1$. Suppose that Assumption 1 holds. Let $p\in(1,\infty)$ and $q\in(1,\infty)$ be such that
$\displaystyle W^{1,p}({T})\hookrightarrow L^{q}({T}),$ (5.1)
that is $1-\frac{d}{p}\geq-\frac{d}{q}$ and $q\leq p$. Then,
$\displaystyle\|\varphi_{h}\|_{L^{q}(\Omega)}\leq C^{dS0(q,p)}|\varphi_{h}|_{p,V_{h}}\quad\forall\varphi_{h}\in V_{h}\in\{V_{h}^{DCCR},V_{h}^{CR},V_{h0}^{CR}\},$ (5.2)
where $C^{dS0(q,p)}$ is a positive constant independent of $h$ and $\varphi_{h}$.
@@
By using the estimate (2.25) for each element, we obtain the following
$\displaystyle\|\varphi_{h}-\Pi_{h}^{0}\varphi_{h}\|_{L^{q}(\Omega)}$ $\leq c|\varphi_{h}|_{W^{1,p}(\mathbb{T}_{h})}.$ (5.4)
Because $\Omega$ is bounded and $q\leq p$, then $L^{p}(\Omega)\hookrightarrow L^{q}(\Omega)$ and
$\displaystyle\|\Pi_{h}^{0}\varphi_{h}\|_{L^{q}(\Omega)}$ $\leq C(\Omega)\|\Pi_{h}^{0}\varphi_{h}\|_{L^{p}(\Omega)}.$ (5.5)
@@
$|\bm{v}_{0}|_{W^{1,p^{\prime}}(\Omega)^{d}}$ $\leq C_{p^{\prime}}\|\xi_{h}\|_{L^{p^{\prime}}(\Omega)}.$
@@
$\leq C(p^{\prime},d,\Omega)\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}\leq c,$
@@
$\displaystyle\|\bm{v}_{0}\|_{W^{1,p^{\prime}}(\Omega)^{d}}$ $\leq C(p^{\prime},\Omega)\|\xi_{h}\|_{L^{p^{\prime}}(\Omega)}\leq C(p^{\prime},\Omega)\left(\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}+\|\bar{\psi}_{h}\|_{L^{p^{\prime}}(\Omega)}\right)$
$\leq C(p^{\prime},d,\Omega)\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}.$
@@
$\displaystyle\|I_{h}^{RT}\bm{v}\|_{L^{p^{\prime}}(\Omega)^{d}}$ $\leq c\|\bm{v}\|_{W^{1,p^{\prime}}(\Omega)^{d}}\leq c\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}\leq c.$
@@
$\displaystyle\left|\langle\Pi_{h}^{0}\varphi_{h},\psi\rangle\right|$ $\leq c|\varphi_{h}|_{W^{1,p}(\mathbb{T}_{h})}+c|\varphi_{h}|_{p,J}\leq c|\varphi_{h}|_{p,V_{h}},$
@@
$\displaystyle\|\Pi_{h}^{0}\varphi_{h}\|_{L^{p}(\Omega)}$ $\leq c|\varphi_{h}|_{p,V_{h}}.$ (5.6)
Gathering the inequalities (5.3), (5.4), (5.5) and (5.6), we conclude the target estimate. ∎
@@ 6 Concluding remarks
However, when $q\mathop{\textgreater}p$, it remains to show the following restricted discrete Sobolev inequality, see [10, Lemma 4]. For that purpose, we imposed a weak elliptic regularity assumption to obtain a discrete Sobolev inequality: Let $p\in[2,\infty)$ if $d=2$, or $p\in[2,6]$ if $d=3$, then $p^{\prime}\in(\frac{2d}{d+2},2]$. We assume that, for any $g\in L^{p^{\prime}}(\Omega)$, the variational problem
@@
holds. In addition to the weak elliptic regularity assumption, we impose that Assumptions 1 and there exists a positive constant $C^{dS0}$ independent of $T$ and $h$ such that

### v2

@@ 4 Lemmata for analysis
Furthermore, the operator $\mathcal{L}$ maps $\mathcal{D}(D)\cap L^{1}_{0}(D)$ into $\mathcal{D}(D)^{d}$ and for each real $r\in(1,\infty)$, is continuous from $L^{r}_{0}(D)$ into $W^{1,r}_{0}(D)^{d}$; there exists a constant $C^{Bog}(d,D,r)$} such that
$\displaystyle|\mathcal{L}f|_{W^{1,r}(D)^{d}}\leq{C^{Bog}(d,D,r)}\|f\|_{L^{r}(D)}\quad\forall f\in L^{r}_{0}(D),$ (4.2)
where $L^{r}_{0}(D):=\{v\in L^{r}(D):\ \int_{D}vdx=0\}$. Here, the constant $C^{Bog}(d,D,r)$ depends only on $r$, $d$, and the geometry of $D$, and is, in particular independent of the mesh $\mathbb{T}_{h}$. Furthermore, if $D$ is star-shaped with respect to a ball $B_{\rho}\subset D$ of radius $\rho$, then $C^{Bog}(d,D,r)$ admits an upper bound in terms of $d$, $r$, and the chunkiness parameter $\eta_{D}:=\frac{\mathop{\mathrm{diam}}(D)}{\rho}$. For more general domains obtained by patching such subdomains (e.g. via a partition-of-unity argument), the dependence of the constant on the geometry may be less explicit.
@@
Existence and $W^{1,r}$ -stability of a right inverse of the divergence on Lipschitz-type domains can be found, for instance, in [4, Theorem 1.4.5], where no explicit tracking of the continuity constant is provided. On domains that are star-shaped with respect to a ball, one may obtain bounds for the stability constant in terms of the chunkiness parameter; see, e.g., [10, Chapter III.3]. Extensions to domains representable as finite unions of such star-shaped subdomains can be achieved via a partition-of-unity argument; see, e.g., [11, Section 5] and references therein. We omit the details.
@@
The identity is established through an elementwise application of integration by parts to $\bm{\tau}_{h}\cdot\nabla_{h}\psi_{h}$ and $\operatorname{div}\bm{\tau}_{h}\,\psi_{h}$, in conjunction with the definitions of the jumps and weighted averages. The standard procedural details are not included here.
@@
Let $p^{\prime}\in[1,\infty)$. For any $\bm{w}\in W^{1,p^{\prime}}(\Omega)^{d}$, we set $\bm{\tau}_{h}:={I}_{h}^{RT}\bm{w}$ in (4.3). Using the definition of $I_{h}^{RT}$ yields (4.4).
@@
Let $p^{\prime}\in(1,\infty)$. For any $\bm{w}\in W^{1,p^{\prime}}(\Omega)^{d}$ and $\psi_{h}\in V_{h}^{DC(1)}$,
$\displaystyle\left|\sum_{F\in\mathcal{F}_{h}^{i}}\int_{F}\{\!\{\bm{w}\}\!\}_{\omega,F}\cdot\bm{n}_{F}\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}ds\right|$ $\leq{C_{0a}(d,p^{\prime})}|\psi_{h}|_{p,J}\|\bm{w}\|_{W^{1,p^{\prime}}(\Omega)^{d}},$ (4.5)
$\left|\sum_{F\in\mathcal{F}_{h}^{\partial}}\int_{F}(\bm{w}\cdot\bm{n}_{F})\Pi_{F}^{0}\psi_{h}ds\right|$ $\leq{C_{0b}(d,p^{\prime})}|\psi_{h}|_{p,J}\|\bm{w}\|_{W^{1,p^{\prime}}(\Omega)^{d}},$ (4.6)
where $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. Here, the constants $C_{0a}(d,p^{\prime})$ and $C_{0b}(d,p^{\prime})$ depend only on $d$ and $p^{\prime}$, and are independent of the mesh size $h$ and of the semi-regular mesh parameter $\gamma_{0}$.
@@
Suppose that $F\in\mathcal{F}_{h}^{i}$ with $F=\partial T_{+}\cap\partial T_{-}$, $T_{+},T_{-}\in\mathbb{T}_{h}$. Using the Hölder’s inequality, the weighted average and the trace inequality (2.1) yields
@@
$\quad\leq{C^{Tr}(d,p^{\prime})}\left(\|\bm{w}_{+}\|_{L^{p^{\prime}}(T_{+})^{d}}+h_{T_{+}}^{\frac{1}{p^{\prime}}}\|\bm{w}_{+}\|_{L^{p^{\prime}}(T_{+})^{d}}^{1-\frac{1}{p^{\prime}}}|\bm{w}_{+}|_{W^{1,p^{\prime}}(T_{+})^{d}}^{\frac{1}{p^{\prime}}}\right)\omega_{T_{+},F}\ell_{T_{+},F}^{-\frac{1}{p^{\prime}}}\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}$
$\quad\quad+{C^{Tr}(d,p^{\prime})}\left(\|\bm{w}_{-}\|_{L^{p^{\prime}}(T_{-})^{d}}+h_{T_{-}}^{\frac{1}{p^{\prime}}}\|\bm{w}_{-}\|_{L^{p^{\prime}}(T_{-})^{d}}^{1-\frac{1}{p^{\prime}}}|\bm{w}_{-}|_{W^{1,p^{\prime}}(T_{-})^{d}}^{\frac{1}{p^{\prime}}}\right)\omega_{T_{-},F}\ell_{T_{-},F}^{-\frac{1}{p^{\prime}}}\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}$
$\quad\leq{C^{Tr}(d,p^{\prime})}\Bigg\{\left(\|\bm{w}_{+}\|_{L^{p^{\prime}}(T_{+})^{d}}+h_{T_{+}}^{\frac{1}{p^{\prime}}}\|\bm{w}_{+}\|_{L^{p^{\prime}}(T_{+})^{d}}^{1-\frac{1}{p^{\prime}}}|\bm{w}_{+}|_{W^{1,p^{\prime}}(T_{+})^{d}}^{\frac{1}{p^{\prime}}}\right)^{p^{\prime}}$
@@
Using the Hölder inequality again and the fact that $h\leq 1$, we have
@@
$\quad\leq{C_{1}(d,p^{\prime})}\sum_{F\in\mathcal{F}_{h}^{i}}\left(\omega_{T_{+},F}^{p}\ell_{T_{+},F}^{-\frac{p}{p^{\prime}}}+\omega_{T_{-},F}^{p}\ell_{T_{-},F}^{-\frac{p}{p^{\prime}}}\right)^{\frac{1}{p}}\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}$
@@
$\quad\leq{C_{2}(d,p^{\prime})}\left(\sum_{F\in\mathcal{F}_{h}^{i}}\left(\omega_{T_{+},F}^{p}\ell_{T_{+},F}^{-\frac{p}{p^{\prime}}}+\omega_{T_{-},F}^{p}\ell_{T_{-},F}^{-\frac{p}{p^{\prime}}}\right)\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}^{p}\right)^{\frac{1}{p}}$
@@
$\quad\leq{C_{0a}(d,p^{\prime})}|\psi_{h}|_{p,J}\|\bm{w}\|_{W^{1,p^{\prime}}(\Omega)^{d}},$
which is the inequality (4.5) together with the weight (2.9) and Jensen inequality, where the constants $C_{1}(d,p^{\prime})$ and $C_{2}(d,p^{\prime})$ depend only on $d$ and $p^{\prime}$, and are independent of the mesh size $h$ and of the semi-regular mesh parameter $\gamma_{0}$. Here, $\mathbb{T}_{F}$ denotes the set of the simplices in $\mathbb{T}_{h}$ that share $F$ as a common face.
By an analogous argument, the estimate (4.6) holds.
@@ 5 Main theorem
Suppose that Assumption 1 holds. Let $p\in(1,\infty)$ and $q\in(1,\infty)$ be such that (2.22), that is $1-\frac{d}{p}\geq-\frac{d}{q}$ and $q\leq p$. Then,
$\displaystyle\|\varphi_{h}\|_{L^{q}(\Omega)}\leq{C^{dS0}(d,\Omega,p,q,\gamma_{0})}|\varphi_{h}|_{p,V_{h}}\quad\forall\varphi_{h}\in V_{h}\in\{V_{h}^{DCCR},V_{h}^{CR},V_{h0}^{CR}\},$ (5.1)
where the constant $C^{dS0}(d,\Omega,p,q,\gamma_{0})$ depends only on $d$, $\Omega$, $p$, $q$ and the semi-regular mesh parameter $\gamma_{0}$, and is independent of $h$ and of the aspect ratios and angles of the simplices.
@@
Because $\Omega$ is bounded and $q\leq p$, by using the estimate (2.23) for each element, we obtain the following
$\displaystyle\|\varphi_{h}-\Pi_{h}^{0}\varphi_{h}\|_{L^{q}(\Omega)}$ $\leq{C_{3}(d,\Omega,p,q)}|\varphi_{h}|_{W^{1,p}(\mathbb{T}_{h})},$ (5.3)
where $C_{3}(d,\Omega,p,q)$ depends only on $d$, $\Omega$, $p$ and $q$, and is independent of $h$ and of the aspect ratios and angles of the simplices. Furthermore, because $\Omega$ is bounded and $q\leq p$, then $L^{p}(\Omega)\hookrightarrow L^{q}(\Omega)$ and
$\displaystyle\|\Pi_{h}^{0}\varphi_{h}\|_{L^{q}(\Omega)}$ $\leq{C_{4}(\Omega,p,q)}\|\Pi_{h}^{0}\varphi_{h}\|_{L^{p}(\Omega)},$ (5.4)
where $C_{4}(\Omega,p,q)$ depends only on $\Omega$, $p$ and $q$, and is independent of $h$ and of the aspect ratios and angles of the simplices.
@@
$|\bm{v}_{0}|_{W^{1,p^{\prime}}(\Omega)^{d}}$ $\leq{C^{Bog}(d,\Omega,p^{\prime})}\|\xi_{h}\|_{L^{p^{\prime}}(\Omega)}.$
@@
$\leq{C_{5}(d,\Omega,p^{\prime})}\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}\leq{C_{5}(d,\Omega,p^{\prime})},$
@@
$\displaystyle\|\bm{v}_{0}\|_{W^{1,p^{\prime}}(\Omega)^{d}}$ $\leq{C_{6}(d,\Omega,p^{\prime})}\|\xi_{h}\|_{L^{p^{\prime}}(\Omega)}\leq{C_{6}(d,\Omega,p^{\prime})}\left(\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}+\|\bar{\psi}_{h}\|_{L^{p^{\prime}}(\Omega)}\right)$
$\leq{C_{7}(d,\Omega,p^{\prime})}\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}.$
Here, $C_{5}(d,\Omega,p^{\prime})$, $C_{6}(d,\Omega,p^{\prime})$ and $C_{7}(d,\Omega,p^{\prime})$ depend only on $d$, $\Omega$ and $p^{\prime}$ and is independent of $h$ and of the aspect ratios and angles of the simplices.
@@
$\displaystyle\|I_{h}^{RT}\bm{v}\|_{L^{p^{\prime}}(\Omega)^{d}}$ $\leq{C^{RTsta}(d,p^{\prime},\gamma_{0})}\|\bm{v}\|_{W^{1,p^{\prime}}(\Omega)^{d}}\leq{C_{8}(d,\Omega,p^{\prime},\gamma_{0})}\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}\leq{C_{8}(d,\Omega,p^{\prime},\gamma_{0})},$
where $C_{8}(d,\Omega,p^{\prime},\gamma_{0})$ depends only on $d$, $\Omega$, $p^{\prime}$ and the semi-regular mesh parameter $\gamma_{0}$, and is independent of $h$ and of the aspect ratios and angles of the simplices.
@@
$\displaystyle\left|\langle\Pi_{h}^{0}\varphi_{h},\psi\rangle\right|$ $\leq{C_{8}(d,\Omega,p^{\prime},\gamma_{0})}|\varphi_{h}|_{W^{1,p}(\mathbb{T}_{h})}+{C_{9}(d,p^{\prime})}|\varphi_{h}|_{p,J}\leq{C_{10}(d,\Omega,p^{\prime},\gamma_{0})}|\varphi_{h}|_{p,V_{h}},$
@@
$\displaystyle\|\Pi_{h}^{0}\varphi_{h}\|_{L^{p}(\Omega)}$ $\leq{C_{10}(d,\Omega,p^{\prime},\gamma_{0})}|\varphi_{h}|_{p,V_{h}}.$ (5.5)
Here, $C_{9}(d,p^{\prime})$ depends only on $d$ and $p^{\prime}$, and $C_{10}(d,\Omega,p^{\prime},\gamma_{0})$ depends only on $d$, $\Omega$, $p^{\prime}$ and the semi-regular mesh parameter $\gamma_{0}$. They are independent of $h$ and of the aspect ratios and angles of the simplices.
Gathering the inequalities (5.2), (5.3), (5.4) and (5.5), we conclude the target estimate.
@@ 6 Concluding remarks
However, when $q\mathop{\textgreater}p$, it remains to show the following restricted discrete Sobolev inequality, see [16, Lemma 4]. For that purpose, we imposed a weak elliptic regularity assumption to obtain a discrete Sobolev inequality: Let $p\in[2,\infty)$ if $d=2$, or $p\in[2,6]$ if $d=3$, then $p^{\prime}\in(\frac{2d}{d+2},2]$. We assume that, for any $g\in L^{p^{\prime}}(\Omega)$, the variational problem
@@
holds. In addition to the weak elliptic regularity assumption, we impose that Assumptions 1 and there exists a positive constant $C^{(p,2)}$ independent of $T$ and $h$ such that